OpenStudy (anonymous):

Will reward with Medal! What equation results from completing the square and then factoring? x^2 - 10x = 46 please help.

OpenStudy (anonymous):

My choices are here ... I'm confused

OpenStudy (mathmale):

have you "completed the square" before? If so, what is the first step?

OpenStudy (anonymous):

Not sure, could you go through some of the steps?

OpenStudy (mathmale):

Look at x^2 - 10x = 46. Re-write it as 1x^2 - 10 x =46 and compare this result to: ax^2 + bx + c. a=? b=?

OpenStudy (anonymous):

?

OpenStudy (mathstudent55):

x^2 - 10x = 46 To complete the square, start with an equation that has x^2, not 2x^2 or 3x^2, etc. In this case, you already have that. Then move the constant term (term with no variable) to the right side. In this case this is already done since you have 46 on the right side.

OpenStudy (anonymous):

Oh so it was already on the right side

OpenStudy (mathstudent55):

The "complete the square" step is next. Take the coefficient of the x term. Take half of it. Then square that. Then add that to both sides.

OpenStudy (anonymous):

apex?

OpenStudy (mathmale):

Johnny: those a, b and c are called coefficients. Specifically, they are the coefficients of the x-squared and x term of your polynomial expression. I asked you to do a comparison, a few lines earlier. a=1 and b =-10. Be sure to spend enough time on this so that you understand and remember it and can identify those coefficients. Again, b=-10. Take half of that, please (divide -10 by 2). What's the result?

OpenStudy (anonymous):

-5

OpenStudy (mathstudent55):

x^2 - 10x + _____ = 46 + _____ The coefficient of the x term is -10. Half of -10 is -5. -5 squared is 25. Add 25 to both sides and place the 25 as shown below (in the blanks above) x^2 - 10x + 25 = 46 + 25

OpenStudy (mathmale):

Good. Johnny: now square that result. In other words, square (-5). (-5)^2=??

OpenStudy (mathstudent55):

Now factor the left side. It is the square of a binomial. Also, add 46 and 25 on the right side. That'll give you your answer.

OpenStudy (anonymous):

It's 71 So my answer would be D

OpenStudy (mathmale):

Right. In summary, Johnny: (-5)^2=25. As mathstudent55 has already done, add 25 to both sides of your equation. You'll get \[x^2 - 10x + 25 = 46 + 25\]

OpenStudy (anonymous):

I believe

OpenStudy (anonymous):

Both were helpful, thank you.

OpenStudy (mathmale):

Please let's go through this problem solution step-by-step (even though I imagine you're anxious to get the "answer." Actually, (-5)^2=25. It's important that you recognize that you now have a perfect square trinomial on the left side. (On the right side you have the sum 69, which is also a perfect square. I'm going to ask you to factor x^2 - 10x +25. This is a crucial skill in algebra. What are the factors?

OpenStudy (mathmale):

@Johnny621 : I'm hoping you'll stick with this until we find the final solution. Sorry to say, it's not 71.

OpenStudy (anonymous):

It worked out

OpenStudy (anonymous):

So how could it not be?

OpenStudy (anonymous):

Maybe just different ways of going about it ...

OpenStudy (mathmale):

Would you mind demonstrating just what you did? I've asked you to factor x^2 - 10 x + 25; would you mind doing that here and now?

OpenStudy (anonymous):

As the other user said above, I took half of -10 then squared it. After that I took the answer and added it to the right side

OpenStudy (anonymous):

46+25 = 71

OpenStudy (mathmale):

Here's what I was looking for: x^2 - 10x + 25 = 71 Factoring the left side produces (x-5)^2 = 71.

OpenStudy (mathmale):

How many answers would there be ? How many x values?

OpenStudy (anonymous):

So you claim that it wasn't 71 but now it is?

OpenStudy (anonymous):

Confusing my sleek-feathered one.

OpenStudy (mathmale):

If (x-5)^2 = 71, what happens when you take the square root of both sides? Johnny, if you don't care to spend the additional time to understand this question thoroughly, you don't have to. Please make up your mind. The answer is NOT 71. Rather, the expression that leads to solutions is (x-5)^2 = 71.

OpenStudy (anonymous):

I wasn't saying that the answer entirely was "71" I was solving the next step that I was told to do. After I had done that step I had figured out the answer was "D".

OpenStudy (anonymous):

Is there always a stick in your bum?

OpenStudy (anonymous):

I understand you're only trying to help and I think that's great, don't get me wrong.

OpenStudy (anonymous):

It's hard to tell what your intention is ...

OpenStudy (anonymous):

You can help someone else if you're still around. I'll probably be back with another question soon though.

OpenStudy (mathmale):

Lost my connection to the 'Net; sorry. I'd suggest you be careful with your language: No matter how young or how old I am or how young or how old you are, "stick in your bum" is highly inappropriate here.

OpenStudy (anonymous):

There are a lot of other phrases I had in mind ... anyway I'll try to think of something more "appropriate" next time.

OpenStudy (mathmale):

I've noticed that you still have not answered several key, mathematical questions that I've asked you. I'm a retired math professor and I have a special fondness for asking students to demonstrate their knowledge instead of just pouring out answers. Unfortunately, if you can't or won't answer, it tells me you may not very well understand what we're doing. So I ask that you please try to answer my questions next time.

OpenStudy (mathmale):

I'd suggest you just move on; work some other math problems. The more practice you get, the better.

OpenStudy (anonymous):

Math isn't my best subject as you may have noticed. I had already got my answer. I knew the answer wasn't "71" and I apologize that you may have thought that was the case. I really don't understand much and I don't want to spend much time trying to. That's my choice as you said earlier. Congratulations on earning the title of retired math professor. I'll be back with another question for sure, try to have a nice day until then!